# Evaluating Decision Optimality of Autonomous Driving via Metamorphic Testing

### v1

@@ 3.2. Approach Overview
Figure 2 provides a high-level depiction of our testing framework Decictor, designed to detect non-optimal decisions using the metamorphic relation. The basic idea underlying our method involves generating a scenario $s^{\prime}$, where the ego vehicle opts for a path that is far from the optimal path in $s$. It can be formulated as an optimization problem aiming to maximize the difference between the driving paths or behaviors of $s$ and $s^{\prime}$:
(2) $\Delta_{s}=\operatorname*{arg\,max}_{\Delta}\mathcal{D}(\tau(s),\tau(s^{\prime})),\text{ s.t., }s^{\prime}=\delta(s,\Delta),$
@@
where $\Delta$ symbolizes the metamorphic relation-compliant perturbation on the scenario $s$, and $\mathcal{D}$ represents a feedback function used to measure the distance or difference between the driving paths or behaviors of the two scenarios. Therefore, we can generate the NoDS from $s$: $s^{\prime}=\delta(s,\Delta_{s})$.
@@
Decictor adopts a search-based method to solve the optimization problem. The process initiates with an initial ODS $\hat{s}$ and aims to output a set of MR failure tests, i.e., NoDSs. We first initialize a population $\mathbf{Q}$ based on $\hat{s}$. Decictor then optimizes this population iteratively until the testing budget $\mathbf{B}$ exceeds. In each iteration, Decictor performs Non-invasive Mutation ($\delta$) to generate the offspring $\mathbf{Q^{\prime}}$ that has the same size as $\mathbf{Q}$. To comply with the constraints of the MR, the mutation strategy ensures that the original optimal path in $s$ remains available in the new scenarios. To determine if the ADS makes optimal driving decisions in these new scenarios, we introduce MR Check ($\epsilon$), a hierarchical approach for measuring equivalence between the initial ODS and these new scenarios. MR Check initially monitors whether a new scenario has completed the task in virtual testing. If the new scenario fails this test, it is identified as a task failure and subsequently not regarded as a NoDS. Conversely, if the new scenario passes the task checking, we collect the observation including the driving path from the simulator. The MR Check then verifies the equivalence of the driving paths between the initial ODS and the new scenario by using a lane-grid measurement. Detection of the MR violation in this step identifies a NoDS. If no violation occurs, Decictor proceeds to select the superior population from the candidates of the old and new populations for the next iteration. This selection process is guided by a Feedback ($\mathcal{D}$) that takes into account the spatial and temporal characteristics, i.e., the driving path and the behaviors, of the ego vehicle’s motion.
4. Approach
Input :  An initial seed ODS $\hat{s}$
Output :  A set of NoDSs $\mathbf{F}_{n}$
@@ 4. Approach
1 $\mathbf{F}_{n}\leftarrow\{\}$, $\mathbf{Q}\leftarrow\{\}$
2 for $i\in\{1,\ldots,N\}$ do
3       $\mathbf{Q}\leftarrow\mathbf{Q}\cup\{\hat{s}\}$
@@
5repeat
@@ 6       $\mathbf{Q}^{\prime}\leftarrow\{\}$
7       for ${s}\in\mathbf{Q}$ do
             $s^{\prime}\leftarrow\delta({s},\Delta)$ // Non-invasive Mutation
8             $r_{vt},O(s^{\prime})\leftarrow\textnormal{{{{VirtualTest}}(}}\textnormal{\emph{s'}}\textnormal{{)}}$
9             if $r_{vt}$ is pass then
                   // MR Check
10                   $r_{mr}\leftarrow\epsilon(\tau(s),\tau(s^{\prime}))$
11                   if $r_{mr}$ is pass then
@@ 14                         $\mathbf{F}_{n}\leftarrow\mathbf{F}_{n}\cup\{s^{\prime}\}$
17
18until Testing budget $B$ exhausted;
return $\mathbf{F}_{n}$
@@
Algorithm 1 presents the main algorithmic procedure of Decictor. Decictor receives a seed ODS $s$ as the input and outputs a set of NoDSs violated the metamorphic relation. Parameters $N$ and $B$ can be adjusted to configure the population size and testing budget, respectively. The algorithm begins by creating an initial population with the given initial ODS $\hat{s}$ (Lines 2-3). In each iteration, the algorithm first generates the offspring by mutating each scenario in the population $\mathbf{Q}$ (Lines 6-7). Each new scenario $s^{\prime}$ is first executed by the virtual testing (Line 8) and verified by the task checking (Line 9). The new scenario passed the task checking is then evaluated for the satisfiability of the MR (Lines 10-14). If the MR is violated, a NoDS is identified and added to the set $\mathbf{F}_{n}$ (Line 14). If no MR violation is detected, the new scenario is added to the offspring $\mathbf{Q^{\prime}}$. The scenario selection process then commences, which involves picking the top $N$ scenarios from the union of $\mathbf{Q}$ and $\mathbf{Q^{\prime}}$ based on their fitness scores (Line 15). The algorithm ends by returning detected NoDSs $\mathbf{F}_{n}$ (Line 17).
@@
In the following sections, we introduce the key components of Decictor: the Non-invasive Mutation ($\delta$), the Metamorphic Relation Check ($\epsilon$), and the Feedback ($\mathcal{D}$).
4.1. Non-invasive Mutation $\delta$
The main challenge of the mutation is to ensure that in the mutated scenario, the optimal path of the ego vehicle from the seed ODS is not affected. However, existing mutation techniques for ADS safety testing typically alter the waypoints of NPC vehicles (Li et al., 2020; Cheng et al., 2023; Haq et al., 2022; Tang et al., 2021c), which could significantly impact the ego vehicle’s optimal path in the original ODS. To address these limitations, we propose a non-invasive mutation that effectively generates mutated scenarios without disrupting the optimal path of the ODS. The non-invasive mutation mainly includes adding new participants within the non-invasive feasible areas and removing existing added participants.
Non-invasive Feasible Area. To guarantee non-invasion, we aim to compute a set of non-invasive feasible areas such that the newly added participant $P_{m}$ will not affect the motion of the ego vehicle and other participants if $P_{m}$ moves around in these non-invasive feasible areas. Note that it is essential that the motions of other participants remain unaffected, as their behaviors could indirectly alter the motion of the ego vehicle.
Given an ODS $s=\{\mathcal{A},\mathcal{P}\}$ and the observation $O(s)=\{s_{0},s_{1},\ldots,s_{k}\}$ with a sampling time step $\Delta t$, we define the non-invasive feasible area regarding to $p\in\{\mathcal{P}\cup P_{0}\}$ for the next timestamp $t+\Delta t$ as $R_{p}(s,t,\Delta t)$, where $P_{0}$ is the ego vehicle.
Figure 3. Illustrative example of feasible area calculation.
@@ 4.1. Non-invasive Mutation $\delta$
Figure 3 illustrates the basic idea about the computation of $R_{p}(s,t,\Delta t)$, taking the ego vehicle as an example. First, the new participant should move in the feasible motion area, denoted as $R(y_{t})$, according to its kinematic constraints (e.g., the maximal speed and steering angle) and the position $y_{t}$ at timestamp $t$, i.e., the sector in Figure 3(a). Second, suppose the motion area of the ego vehicle $P_{0}$ between $[t,t+\Delta t]$ is $R^{P_{0}}(s_{t},s_{t+\Delta t})$, e.g., the dashed light orange area in Figure 3(b). $R^{P_{0}}(s_{t},s_{t+\Delta t})$ can be determined as $P_{0}$’s motion area from $s_{t}$ to $s_{t+\Delta t}$ in the ODS $s$. Clearly, the new participant should not move into $R^{P_{0}}(s_{t},s_{t+\Delta t})$ during $[t,t+\Delta t]$. Therefore, we have
(3) $R_{P_{0}}(s,t,\Delta t)=R(y_{t})\setminus R^{P_{0}}(s_{t},s_{t+\Delta t}),$
i.e., the light green area in Figure 3(b).
Consequently, the non-invasive feasible area for the newly added participant at timestamp $t+\Delta t$ is:
(4) $R(s,t,\Delta t)=\cap_{p\in\{\mathcal{P}\cup P_{0}\}}R_{p}(s,t,\Delta t)$
where $R_{p}(s,t,\Delta t)$ is the non-invasive feasible area regarding to participant $p$, and $P_{0}$ is the ego vehicle.
@@
It is worth noting that the time step $\Delta t$ will affect the computation of the non-invasive area and the detection of NoDSs significantly. As $\Delta t$ increases, the non-invasive area becomes more conservative, leading to fewer detected NoDSs.
Input :  Participants in the initial ODS $\hat{s}$: $\mathcal{P}_{\hat{s}}$
Participants in the current ODS $s$: $\mathcal{P}_{s}$
Observation of $s$: $O(s)=\{s_{0},s_{1},\ldots,s_{k}\}$
@@ 2       $P_{m}\leftarrow\{\}$
3       $y_{0}\leftarrow$ a collision-free waypoint with zero speed
@@
5             $R(s,t,\Delta t)\leftarrow$ NonInvasiveArea($y_{t},O(s),\Delta t$)
6             if $R(s,t,\Delta t)$ is $\emptyset$ then
@@ 5             $R(s,t,\Delta t)\leftarrow$ NonInvasiveArea($y_{t},O(s),\Delta t$)
8            $y_{t+\Delta t}\leftarrow$ Sample($R(s,t,\Delta t)$)
@@ 9             $P_{m}\leftarrow P_{m}\cup\{y_{t+\Delta t}\}$
10      $\mathcal{P}_{s^{\prime}}\leftarrow\mathcal{P}_{s}\cup\{{P_{m}}\}$
@@
12Function Removing():
13       $P_{m}\leftarrow$ Sample($\mathcal{P}_{s}\setminus\mathcal{P}_{\hat{s}}$)
14       $\mathcal{P}_{s^{\prime}}\leftarrow\{\mathcal{P}_{s}\setminus P_{m}\}$
@@
Mutation Operations. Algorithm 2 outlines the specific mutation operations: Adding (Lines 1-11) and Removing (Lines 12-15). The non-invasive mutation takes as input $\mathcal{P}_{\hat{s}}$ (the participants in the initial ODS $\hat{s}$), $\mathcal{P}_{s}$ (the participants in the current ODS $s$) and $O(s)$ (the observation of $s$). The observation consists of a sequence of scenes $\{s_{0},\ldots,s_{k}\}$ with an equal time step $\Delta t$. In each iteration, the non-invasive mutation randomly chooses one operation to change participants in the scenario $s$, resulting in a new scenario $s^{\prime}$.
@@
Adding. Adding operation aims to introduce complexity to the scenario by adding a new participant to $s$ without influencing the optimal path. In detail, this operation initializes an empty waypoint set $P_{m}$ and a collision-free waypoint $y_{0}$ (Lines 2-3). Waypoints for the added participant are iteratively generated (Lines 4-9), where each iteration involves calculating the non-invasive area (Line 5), sampling a non-invasive waypoint (Line 6-8), and adding it to $P_{m}$ (Line 9). The process concludes by merging the resultant set with the original participants $\mathcal{P}_{s}$ to form $\mathcal{P}_{s^{\prime}}$ (Line 10-11).
@@
Removing. Continuously adding new participants to the scenario will introduce too many obstacles, resulting in motion task failures. Therefore, we introduce the removing operator in this paper. Note that only the participants in $\mathcal{P}_{s}\setminus\mathcal{P}_{\hat{s}}$, i.e., the participants newly added to the initial ODS, can be removed, while the participants in the initial ODS  will remain unchanged in all mutated scenarios. The reason is that the removal of the participants in the initial ODS may affect the optimal path of the scenarios. Specifically, this operation removes a participant in $\mathcal{P}_{s}\setminus\mathcal{P}_{\hat{s}}$ randomly (Lines 13-14) and returns the mutated set (Line 15).
4.2. MR Check $\epsilon$
Given an ODS $s$ and a mutated scenario $s^{\prime}$, we require a criterion $\epsilon$ to determine whether $s^{\prime}$ remains an ODS. A direct approach might be checking if the driving path in $s^{\prime}$ (i.e., $\tau(s^{\prime})$) and the original path of $s$ (i.e., $\tau(s)$) are the same. However, it can be overly strict, as $\tau(s^{\prime})$ can still be considered optimal even if it slightly deviates from the original path $\tau(s)$ (e.g., the path in Figure 4(b) can still be regarded as an optimal one). To address this, we propose an abstraction-based method to quantify the similarity between the two driving paths.

### v3

@@ III-B Approach Overview
Fig. 2 provides a high-level depiction of our testing framework Decictor, designed to detect non-optimal PPDs. The basic idea underlying our method involves generating a scenario $s^{\prime}$, where the ego vehicle opts for a path far from the optimal one in the ODS $s$ under the non-invasive mutation $\delta$. It can be formulated as an optimization problem aiming to maximize the difference between the driving paths of $s$ and $s^{\prime}$:
$\Delta_{s}=\operatorname*{arg\,max}_{\Delta}\mathcal{D}(\tau(s),\tau(s^{\prime})),\text{ where }s^{\prime}=\delta(s,\Delta)$ (2)
@@
where $\Delta$ symbolizes the non-invasive perturbations on the scenario $s$, and $\mathcal{D}$ is a function to measure the distance or difference between the driving paths of the two scenarios.
@@
Decictor adopts a search-based method to solve the optimization problem. The process initiates with a seed ODS $s^{*}$ and aims to output a set of violation tests, i.e., NoDSs, considering the optimal path in $s^{*}$. We first initialize a population $\mathbf{Q}$ based on ${s^{*}}$. Decictor then optimizes this population iteratively until the testing budget $B$ exceeds. In each iteration, Decictor performs Non-invasive Mutation ($\delta$) to generate the offspring $\mathbf{Q^{\prime}}$ that has the same size as $\mathbf{Q}$. The mutation strategy ensures that the original optimal path in $s^{*}$ remains optimal in the new scenarios. To determine if the ADS makes the optimal PPDs in these new scenarios, we introduce Consistency Check ($\epsilon$), a hierarchical approach for measuring the equivalence with regard to PPD optimality between the seed ODS and these new scenarios. Consistency Check initially monitors whether a new scenario completes the task during simulation execution. If the modified scenario results in a task failure, such as collisions, it is not considered as a NoDS of interest. This is because such failures can already be detected by existing AV testing tools that utilize timeout or collision detection oracles. Conversely, if the new scenario passes the task checking, we collect the observation from the simulator. The Consistency Check then verifies the equivalence of the driving paths between the seed ODS and the new scenario using a lane-grid measurement. Detection of the consistency violation in this step identifies a NoDS. If no violation occurs, Decictor proceeds to select the superior population from the candidates of the old and new populations for the next iteration. This selection process is guided by the Feedback ($\mathcal{D}$) that takes into account the spatial and temporal characteristics, i.e., the driving path and the behaviors, of the ego vehicle’s motion.
IV Approach
1
Input :  The seed ODS: $s^{*}$
Output :  The set of NoDSs: $\mathbf{F}_{n}$
@@ IV Approach
2
[…]
@@
5for $i\in\{1,\ldots,N\}$ do
6       $\mathbf{Q}\leftarrow\mathbf{Q}\cup\{{s}^{*}\}$
7
8
9repeat
@@ 10       $\mathbf{Q}^{\prime}\leftarrow\{\}$
11       for $\hat{s}\in\mathbf{Q}$ do
             $s^{\prime}\leftarrow\delta(\hat{s},\Delta)$ // Non-invasive Mutation
12             $r_{tc},O(s^{\prime})\leftarrow\textnormal{{{{SimExecution}}(}}\textnormal{\emph{$s^{\prime}$}}\textnormal{{)}}$
13
14            if $r_{tc}$ is passed then
                   // Consistency Check
15                   $r_{cc}\leftarrow\epsilon(\tau(s^{*}),\tau(s^{\prime}))$
16                   if $r_{cc}$ is passed then
@@ 19                         $\mathbf{F}_{n}\leftarrow\mathbf{F}_{n}\cup\{s^{\prime}\}$
22
23until Testing budget $B$ exhausted;
24return $\mathbf{F}_{n}$
@@
Algorithm 1 presents the main algorithmic procedure of Decictor. Decictor receives a seed ODS $s^{*}$ as the input and outputs a set of NoDSs. Parameters $N$ and $B$ can be adjusted to configure the population size and testing budget, respectively. The algorithm begins by creating an initial population with the given seed ODS $s^{*}$ (Lines 2-3). In each iteration, the algorithm first generates the offspring by mutating each scenario in the population $\mathbf{Q}$ (Lines 6-7). Each new scenario $s^{\prime}$ is first executed in the simulation platform consisting of a simulator and the ADS under test (Line 8) and verified by the task checking $r_{tc}$ (Line 9). The new scenario passed the task checking is then evaluated for PPD consistency (Lines 10-14). If violated, a NoDS is identified and added to the set $\mathbf{F}_{n}$ (Line 14). If no violation is detected, the new scenario is added to the offspring $\mathbf{Q^{\prime}}$. The scenario selection picks the top $N$ scenarios from the union of $\mathbf{Q}$ and $\mathbf{Q^{\prime}}$ based on their fitness scores (Line 15). The algorithm ends by returning detected NoDSs $\mathbf{F}_{n}$ (Line 17).
@@
In the following sections, we introduce the key components of Decictor: the Non-invasive Mutation ($\delta$), the Consistency Check ($\epsilon$), and the Feedback ($\mathcal{D}$).
IV-A Non-invasive Mutation $\delta$
The main challenge of the mutation is to ensure that in the mutated scenario, the original optimal path of the ego vehicle in the seed ODS is not affected. Following existing mutation techniques for ADS safety testing [8, 9, 10, 33], we only alter the waypoints of NPC vehicles or other participants, leaving the ego vehicle unchanged. However, the primary challenge is that these mutations could significantly impact the ego vehicle’s original optimal path. To mitigate this challenge, we introduce a conservative mutation approach, namely the non-invasive mutation strategy, designed to create new scenarios with less impact on the original optimal path. While it may be theoretically difficult, or even impossible, to ensure that changes do not affect the ego vehicle’s original optimal path, our method adopts a conservative approach. This involves calculating “safe zones” for mutation, where adjustments to participants’ behaviors are unlikely to influence the original optimal path. This concept is akin to existing adversarial attack methodologies, where the difference between the new and original inputs is limited within an $L_{P}$ norm boundary [32]. Such constraints are intended to preserve the semantic integrity of the input, but a theoretical guarantee of no impact is challenging to establish [21, 34].
The non-invasive mutation mainly includes adding new participants within the non-invasive feasible areas, removing existing added participants and changing existing participants.
Non-invasive Feasible Area. To minimize the impact of the mutations on the original optimal path $\pi(s^{*})$, for any ODS $\hat{s}=\{\mathcal{A},\mathcal{P}\}$ derived from the seed ODS $s^{*}$, we initially calculate a set of non-invasive feasible areas. These areas are designated so that the introduction of a new participant $P_{m}$ to $\hat{s}$ is unlikely to influence the optimal path $\pi(s^{*})$ and other participants $\mathcal{P}$ in $\hat{s}$ if $P_{m}$ navigates within these non-invasive zones. Note that reducing the influence on the motions of other participants is also important, as their behaviors could indirectly influence the driving path of the ego vehicle.
The determination of the feasible area at any given moment ($t$) must account for the effects of both time and space. Specifically, it involves ensuring that the ego vehicle’s original optimal path $\pi(s^{*})$ and the paths of $\mathcal{P}$ remain unaffected during a certain time frame ($\Delta t$) surrounding that moment. Suppose the current observation is $O(\hat{s})=\{\hat{o}_{0},\hat{o}_{1},\ldots,\hat{o}_{k}\}$ with a sampling time step $\Delta t$, we define the non-invasive feasible area regarding to the object $p\in\{\mathcal{P}\cup P_{0}\}$ during $[t,t+\Delta t]$ as $R_{p}(\hat{s},t,\Delta t)$, where $P_{0}$ is the ego vehicle.
Figure 3: Illustrative example of feasible area calculation.
@@ IV-A Non-invasive Mutation $\delta$
Fig. 3 illustrates the basic idea about the computation of $R_{p}(\hat{s},t,\Delta t)$, taking the ego vehicle as an example. We first estimate the feasible motion area of the new participant, denoted as $R(y_{t})$. Specifically, given the motion constraints (i.e., speed and steering ranges) and the position $y_{t}$ at timestamp $t$, we can calculate a sector range in Fig. 3(a) according to the kinematic model [35]. Second, we estimate the motion area of the ego vehicle $P_{0}$ between $[t,t+\Delta t]$ based on the original optimal path $\pi(s^{*})$, which is denoted as $R^{P_{0}}(\pi(s^{*}),\Delta t)$, e.g., the dashed light red area in Fig. 3(b). Note that the motion area of the ego vehicle is directly calculated from the real movement recorded in the scenario observations, which can be regarded as a rectangle between $t$ and $t+\Delta_{t}$. Intuitively, the new participant should not move into $R^{P_{0}}(\pi(s^{*}),\Delta t)$ during the time frame $[t,t+\Delta t]$ so as not to affect the optimal path of the ego vehicle.
Therefore, we can calculate the non-invasive feasible area (e.g., the light yellow area in Fig. 3(b)) as:
$R_{P_{0}}(\hat{s},t,\Delta t)=R(y_{t})\setminus R^{P_{0}}(\pi(s^{*}),\Delta t),$ (3)
Similarly, we can calculate the non-invasive feasible areas $R_{p}(\hat{s},t,\Delta t)$ with respect to any participant $p\in\mathcal{P}$ based on its driving path $\pi^{p}(\hat{s})$. The inference of the non-invasive feasible area at timestamp $t+\Delta t$ should consider all newly added participants, which is calculated as:
$R(\hat{s},t,\Delta t)=\cap_{p\in\{\mathcal{P}\cup P_{0}\}}R_{p}(\hat{s},t,\Delta t).$ (4)
@@
Note that the time step $\Delta t$ will affect the computation of the non-invasive area. As $\Delta t$ increases, the calculation of the non-invasive area becomes more conservative as the new participant does not affect the ego behavior in longer time.
1
Input : Participants in the seed ODS $s^{*}$: $\mathcal{P}_{s^{*}}$
The ego vehicle’s initial optimal path: $\pi(s^{*})$
Participants in the current ODS $\hat{s}$: $\mathcal{P}_{\hat{s}}$
Observation of $\hat{s}$: $O(\hat{s})=\{\hat{o}_{0},\hat{o}_{1},\ldots,\hat{o}_{k}\}$
@@ 4       $P_{m}\leftarrow\{\}$
5       $y_{0}\leftarrow$ a collision-free waypoint
@@
7             $R(\hat{s},t,\Delta t)\leftarrow$ NonInvasiveArea($y_{t},\pi(s^{*}),O(\hat{s}),\Delta t$)
8             if $R(\hat{s},t,\Delta t)$ is $\emptyset$ then
@@ 7             $R(\hat{s},t,\Delta t)\leftarrow$ NonInvasiveArea($y_{t},\pi(s^{*}),O(\hat{s}),\Delta t$)
10
11            $y_{t+\Delta t}\leftarrow$ Sample($R(\hat{s},t,\Delta t)$)
@@ 12             $P_{m}\leftarrow P_{m}\cup\{y_{t+\Delta t}\}$
13      $\mathcal{P}_{s^{\prime}}\leftarrow\mathcal{P}_{\hat{s}}\cup\{{P_{m}}\}$
@@
15
16Function Removing():
17       $P_{m}\leftarrow$ Sample($\mathcal{P}_{\hat{s}}\setminus\mathcal{P}_{s^{*}}$)
18       $\mathcal{P}_{s^{\prime}}\leftarrow\{\mathcal{P}_{\hat{s}}\setminus P_{m}\}$
@@
20
21 Function Changing():
22       $\mathcal{P}_{s^{\prime}}\leftarrow\textnormal{{{{Removing}}(}}\textnormal{\emph{}}\textnormal{{)}}$
23       $O(\hat{s})\leftarrow O(\hat{s})\setminus\mathbf{y}^{\{\mathcal{P}_{\hat{s}}\setminus\mathcal{P}_{s^{\prime}}\}}$
24       $\mathcal{P}_{\hat{s}}\leftarrow\mathcal{P}_{s^{\prime}}$
25       $\mathcal{P}_{s^{\prime}}\leftarrow\textnormal{{{{Adding}}(}}\textnormal{\emph{}}\textnormal{{)}}$
26       return $\mathcal{P}_{s^{\prime}}$
27
@@ 25       $\mathcal{P}_{s^{\prime}}\leftarrow\textnormal{{{{Adding}}(}}\textnormal{\emph{}}\textnormal{{)}}$
Mutation Operations. Algorithm 2 outlines the specific mutation operations: Adding (Lines 1-11), Removing (Lines 12-15) and Changing (Line 16-21). The non-invasive mutation takes as input $\mathcal{P}_{s^{*}}$ (the participants in the seed ODS $s^{*}$), $\mathcal{P}_{\hat{s}}$ (the participants in the current ODS $\hat{s}$) and $O(\hat{s})$ (the observation of $\hat{s}$). The observation consists of a sequence of scenes $\{\hat{s}_{0},\ldots,\hat{s}_{k}\}$ with an equal time step $\Delta t$. In each iteration, the non-invasive mutation randomly chooses one operation to change participants in the scenario $\hat{s}$, resulting in a new scenario $s^{\prime}$.
@@
Adding. Adding operation aims to introduce complexity to the scenario by adding a new participant to $\hat{s}$ that is less likely to influence the optimal path. In detail, this operation initializes an empty waypoint set $P_{m}$ and an initial collision-free waypoint $y_{0}$ (Lines 2-3). Waypoints for the added participant are iteratively generated (Lines 4-9), where each iteration involves calculating the non-invasive area (Line 5), sampling a non-invasive waypoint (Line 6-8), and adding it to $P_{m}$ (Line 9). The process concludes by merging the resultant set with the original participants $\mathcal{P}_{\hat{s}}$ to form $\mathcal{P}_{s^{\prime}}$ (Line 10-11).
@@
Removing. Continuously adding new participants to the scenario will introduce too many obstacles, resulting in motion task failures. Thus, we introduce the removing operator. Note that only the participants in $\mathcal{P}_{\hat{s}}\setminus\mathcal{P}_{s^{*}}$, i.e., the participants newly added to the seed ODS $s^{*}$, can be removed, while the participants in the seed ODS will remain unchanged in all mutated scenarios. Because the removal of the participants in the seed ODS may affect the optimal path of the scenarios. Specifically, this operation removes a participant in $\mathcal{P}_{\hat{s}}\setminus\mathcal{P}_{s^{*}}$ randomly (Lines 13-14) and returns the mutated set (Line 15).
Changing. The changing operation aims to modify the scenario $\hat{s}$ by replacing an existing participant in $\mathcal{P}_{\hat{s}}$. This operation involves first removing an existing object from $\mathcal{P}{\hat{s}}$ (Line 17) and then adding a new one (Line 20). Specifically, we remove the observation of the removed participant $\mathbf{y}^{\{\mathcal{P}_{\hat{s}}}$ from $O(\hat{s})$ (Lines 18-19) to ensure that the calculation of the non-invasive feasible area does not include this participant.
IV-B Consistency Check $\epsilon$